OpenStudy (anonymous):

If x = 2 for the equation y = 5x -3, what is the ordered pair (x, y)? A. (2, 7) B. (7, 2) C. (4, 2) D. (2, 4)

OpenStudy (blurbendy):

Put 2 in for in the equation y = 5x -3 your ordered pair for y will be what y equals after you put 2 in

OpenStudy (anonymous):

set x into the equation and solve for y

OpenStudy (anonymous):

y = 5(2)-3 y = 10-3 y = 7 so the ordered pair would be (2,7)

OpenStudy (anonymous):

is the answer c because i just want to check it

OpenStudy (anonymous):

oooooo

OpenStudy (anonymous):

it is a

OpenStudy (anonymous):

do u understand who i did that

OpenStudy (anonymous):

how do i know if 2.7 would2 be first and not 7,2 and 7 not be first

OpenStudy (anonymous):

because the problem says that 2 is x, and we solved for y. An ordered pair is in the form of (x,y) so scince 2=x 2 must be first and 7 second, understand?

OpenStudy (anonymous):

@donavan u there